How many atoms of helium gas fill a spherical balloon of diameter 30.0 cm at 20.08 oC and 1.00atm? What is the average kinetic energy of the helium?
1
Expert's answer
2021-02-06T16:42:11-0500

"V=(4/3)\pi r^3=(4/3)\cdot3.14\cdot 0.3^3=0.113(m^3)"


"pV=\nu RT\to pV=\frac{N}{N_A} RT\to"


"N=\frac{pVN_A}{RT}=\frac{101325\cdot 0.113\cdot 6.02\cdot10^{23}}{8.31\cdot 293.08}=2.83\cdot10^{24}" atoms. Answer


The average kinetic energy of motion of a helium molecule


"E_{avg}=\frac{3}{2}kT=\frac{3}{2}\cdot 1.38\cdot 10^{-23}\cdot 293.08=607\cdot10^{-23}\ (J)"